# Tight MIP Formulations for Optimal Operation and Investment of Storage Including Reserves

### v2

@@ Abstract
Fast and accurate large-scale energy system models are needed to investigate the potential of storage to complement the fluctuating energy production of renewable energy systems. However, standard Mixed-Integer Programming (MIP) models that describe optimal investment and operation of these storage units, including the optional capacity to provide up/down reserves, do not scale well. To improve scalability, the integrality constraints are often relaxed, resulting in Linear Programming (LP) relaxations that allow simultaneous charging and discharging, while this is not feasible in practice. To address this, we derive the convex hull of the solutions for the optimal operation of storage for one time period, as well as for problems including investments and reserves, guaranteeing that no tighter MIP formulation or better LP approximation exists for one time period. When incorporating this convex hull into a multi-period formulation and including it in large-scale energy system models, the improved LP relaxations can better prevent simultaneous charging and discharging, and the tighter MIP can speed up the solving time. We demonstrate this with illustrative case studies of a unit commitment problem and a transmission expansion planning problem.
Index Terms:
Energy storage systems, mixed-integer linear programming (MIP), linear programming (LP), convex hull, tight formulation, optimal investments, reserves.
@@ I-A Motivation
To investigate the optimal investment and operation decisions of storage, we need to formulate realistic models for devices capable of storing energy, such as pumped hydro storage, electric vehicles [1], thermal storage, and some forms of demand response [2]. In some cases, we also want to incorporate reserves, which enable more flexibility by representing the possible increase/decrease to the scheduled amount of charged/discharged energy[3]. Mixed-Integer Programming (MIP) allows us to correctly model the optimal operation of storage, using both binary and continuous variables. Binary variables are required to correctly model reserves and to ensure that the storage cannot charge and discharge simultaneously [4, 5]. Constraints in MIP models can describe the characteristics and capabilities of storage units, such as their minimum and maximum state-of-charge, and their (dis)charging capacity.
However, there are two big problems that arise when using these storage models to solve large-scale problems. The first is that computational problems arise when MIP storage models are included in large-scale energy system optimization models. MIP models generally do not scale well, since existing algorithms for finding the optimal solution to MIP models run in exponential time (in the worst case). A common practice to reduce the computation time of these large energy models is to use a decomposition technique, or to relax the integrality restrictions in the MIP formulations. In either case, the Linear Programming (LP) relaxation of the storage model is solved, which can be solved much faster in practice. However, the problem with these LP models for storage operation is that their solutions can enforce simultaneous charging and discharging, although this cannot be executed in practice. To address both of these challenges, we improve the LP approximation by tightening the MIP models for optimal storage investment and operation, including reserves. When incorporated in a large-scale model, these tighter MIP formulations can speed up the solving time. When incorporated into a large-scale LP model, the improved LP relaxations can better prevent simultaneous charging and discharging.
@@ I-B Background & Literature Review
Charging and discharging at the same time essentially increases losses, which commonly increases costs[6]. Therefore, the optimal solution to the LP relaxation usually does not favor this, but it is beneficial in some cases, such as in the event of negative prices. However, simultaneous charging and discharging can usually not be executed in practice. There are exceptions, such as composite battery systems[7] or other specific cases with multiple and/or fast storage units. Here, simultaneous charging and discharging might actually be executable by charging a fraction of the time period or storage units, and discharging another fraction[8]. Still, this is not executable when the storage unit is already at maximum capacity, and it is typically not possible in practice at all for generic storage units. Therefore, solutions in which this occurs are unwanted and can lead to large errors.
@@
Some examples of situations in which simultaneous charging and discharging in the model solutions can lead to large errors are the following. Arroyo et al. [6] present a transmission expansion planning case study, where the optimal investment plan is actually not feasible in practice, because simultaneous charging and discharging is needed to execute it. Furthermore, immediate load recovery cannot be guaranteed in demand response models with load shifting when simultaneous up and down shifting of load is allowed [9]. Lastly, significant errors can occur when decomposition algorithms are used to solve large MIP problems faster [5]. In these decomposition frameworks, the LP relaxation (of the operations/second stage) is solved many times, thus allowing simultaneous charging and discharging and resulting in errors.
@@
Several attempts to prevent simultaneous charging and discharging in convex problems have been made, such as including pre-contingency operating costs in the objective function [10] or adding conditions for the roundtrip efficiency [11]. However, Arroyo et al. [6] pointed out that it can still occur and illustrated this in two case studies: a unit commitment problem and a transmission expansion planning problem. Thus, an improved LP-approximation of storage models that can better prevent simultaneous charging and discharging would be a big step toward more useful and better scalable storage models.
@@
An improved LP relaxation can also speed up the solving time of MIP storage models. The computational performance of MIP models is highly dependent on their tightness and their size (in terms of the number of constraints and variables)[12]. We say that an MIP model is tight if its LP relaxation describes the convex hull of its solutions. We say that MIP model $A$ is tighter than another MIP model $B$ if the LP relaxation of MIP model $A$ is contained inside the LP relaxation of MIP model $B$, and it is strictly smaller than the other (so $\text{LP}(A)\subsetneq\text{LP}(B)$). Typically, the tightness of MIP models can be improved by adding cuts (additional constraints). However, this also increases the size of the model, which in turn could negatively impact the computational performance.
@@
In related fields, such as unit commitment modeling, much research has been done on finding tighter MIP models[13, 12, 14]. These works take into account the important trade-off between model tightness and model size. For example, Damci-Kurt et al.[15] first consider the unit commitment problem with ramping constraints for a single generator in just two time periods. They give a complete linear description of the convex hull of its feasible solutions. This formulation can be incorporated into a multi-generator and multi-period formulation, resulting in a formulation that is tighter than the previously used ramping models, but of similar size. They then consider the multi-period case, and find exponential classes of facet-defining inequalities. Though these improve the tightness of the multi-period problem formulation, they heavily increase the size of the model. It is only beneficial for the solving time if these constraints are added to the formulation using a separation algorithm, when needed, instead of a priori.
@@
For a basic storage operation MIP model, Pozo [16] recently obtained two different convex hull formulations: a vertex representation and a hyperplane representation. Though not clearly stated in their paper, their formulations describe the convex hull of the problem for only one time period. Additional constraints would be needed to describe the convex hull for a multi-period problem, resulting in a model that likely grows exponentially in size w.r.t. the number of time periods. However, the single-period convex hull formulation can be incorporated into a multi-period formulation, which is then tighter than the commonly used formulation. They demonstrate that incorporating this multi-period LP formulation in a multi-period Set-Point Tracking problem prevents simultaneous charging and discharging more often than the LP relaxation of the commonly used MIP model.
@@
However, there are several downsides to the LP formulation obtained by Pozo[16]. First, we observe that their formulation contains several redundant constraints. Half of the constraints can be removed, which decreases the size of the model. Second, their convex hull formulation does not contain the original binary variable. Therefore, their multi-period LP approximation cannot be used to solve the original multi-period MIP storage operation problem. Last, Pozo does not consider reserves and investment decisions in its model.
@@ I-C Contributions
In summary, the main contributions of this paper are the following: 1) We provide a convex hull formulation for optimal storage operation for one period, which consists of fewer constraints than the formulation found by Pozo[16]. We also provide the corresponding tight MIP model and include it in a multi-period formulation, which can be used to solve the original multi-period storage problem, which is not possible with the LP model by Pozo[16]. 2) We provide tighter MIP models for extended problems that include reserves and investment decisions. We prove that their LP-relaxations provide the convex hull of the feasible solutions to the MIP models for one time period, which guarantees that no tighter MIP or better LP approximation for one time period exists. 3) When incorporated in any type of multi-period energy investment/operation system model, we demonstrate that the LP relaxations of the tighter MIP formulations can decrease the frequency of occurrences of simultaneous charging and discharging. 4) When solving a large-scale multi-period MIP model, we demonstrate that the tighter storage MIP formulations can positively affect the solving time.
The improved MIP formulations and LP relaxations can be used for many different types of planning problems that include storage, as well as for transmission lines. Furthermore, the improved LP approximation can reduce errors when using decomposition frameworks, as explained earlier.
@@ II-A Modeling Storage Operation Problems
Let us consider a storage unit with $\underline{E}=$ 5MWh, $\overline{E}=$50MWh, $\overline{P^{\mathrm{C}}}=\overline{P^{\mathrm{D}}}=$10MW, and $\eta^{\mathrm{C}}=\eta^{\mathrm{D}}=$ 0.5 (for simple computational purposes). Now suppose the unit is already at maximum capacity in a certain time period $t^{*}-1$, so $e_{t^{*}-1}=50$. In the event of negative prices, or some situation where the surrounding energy system needs to get rid of energy, simultaneous charging and discharging might be beneficial, as explained in the introduction. The BO-LP model allows this, so it may output an optimal solutions such as $(e_{t^{*}-1},p^{\mathrm{C}}_{t^{*}},p^{\mathrm{D}}_{t^{*}},\delta_{t^{*}})=(50,8,2,0.8)$ for the next time period $t^{*}$. The state of charge at the end of period $t^{*}$ is then $e_{t^{*}}=e_{t^{*}-1}+\eta^{\mathrm{C}}p^{\mathrm{C}}_{t^{*}}\Delta-\frac{1}{\eta^{\mathrm{D}}}p^{\mathrm{D}}_{t^{*}}\Delta=50+0.5\cdot 8\cdot 1-2\cdot 2\cdot 1$ $=50+4-4=50$. So the model allows a solution where it charges as much as it discharges, keeping the battery at the same state of charge, and satisfying all the constraints. However, this can never be executed in practice. Thus, we want to find and LP model that prevents this.
@@ II-B Modeling Storage Operation Problems Including Reserves
As explained earlier, storage can be used to balance the power grid, which is particularly challenging when more renewable energy systems are implemented. Operating reserves, specifically spinning reserves, are useful in this scenario [3]. The up/down reserve provided by a unit is generally defined as the available capacity that the unit can output on top of its scheduled amount in each time period [17]. For storage units specifically, this means that if the storage unit is scheduled to discharge energy, up reserves can be provided by discharging more energy than scheduled, and down reserves by discharging less (or even charging energy). When charging, up reserves can be provided by charging less (or even discharging), and down reserves by charging more.
@@
$e_{t},\ p^{\mathrm{C}}_{t},\ p^{\mathrm{D}}_{t},\ r^{\mathrm{C}+}_{t},\ r^{\mathrm{C}-}_{t},\ r^{\mathrm{D}+}_{t},\ r^{\mathrm{D}-}_{t}\in\mathbb{R}_{\geq 0}\quad$ $\forall t\in T$ (2i)
@@
A common way to model storage operation reserves with an MIP formulation is by adapting the constraints from the BO-MIP model, resulting in the Basic Operation incl. Reserves MIP model (BOR-MIP model)[18]. Here, (2a) and (2b) replace (1b), since the reserves affect the bounds of the energy storage level and vice versa. Constraint (2c) replaces (1c), bounding the amount of energy that can be charged. Constraint (2c) also bounds the down reserves $r^{\mathrm{C}-}_{t}$ that can be realized by charging more than planned, and (2e) bounds the up reserves $r^{\mathrm{C}+}_{t}$, which can be realized by charging less energy than the planned $p^{\mathrm{C}}_{t}$. Similarly, (2d) replaces (1d), and together with (2f) imposes bounds on $p^{\mathrm{D}}_{t}$, $r^{\mathrm{D}+}_{t}$ and $r^{\mathrm{D}-}_{t}$. Decision variables $r^{-}_{t}$ and $r^{+}_{t}$ then represent the total amount of up and down reserves that can be provided in time period $t$. These total reserves are obtained in (2g) and (2h) and bounded by some value $R^{+}$ and $R^{-}$, respectively.
In this paper, we obtain the convex hull of the BOR-MIP model for one time period. We want to point out that there is an issue with modeling reserves using the BOR-MIP model, namely that it does not fully exploit the flexibility of a storage unit. It only allows limited reserves. We explain this further in Appendix -C. Here we also point out issues with the formulation presented by Momber et al.[19]. This formulation adapts the constraints such that the model does fully exploit the flexibility of the storage unit, but it might not ensure that there is enough capacity to provide the reserves. For those still interested in using the formulation by Momber et al., we provide tight formulations in the online companion[20, Section 4].

### v3

@@ Abstract
Fast and accurate large-scale energy system models are needed to investigate the potential of storage to complement the fluctuating energy production of renewable energy systems. However, standard Mixed-Integer Programming (MIP) models that describe optimal investment and operation of these storage units, including the optional capacity to provide up/down reserves, do not scale well. To improve scalability, the integrality constraints are often relaxed, resulting in Linear Programming (LP) relaxations that allow simultaneous charging and discharging, while this is not feasible in practice. To address this, we derive the convex hull of the solutions for the optimal operation of storage for one time period, as well as for problems including investments and reserves, guaranteeing that no tighter MIP formulation or better LP approximation exists for one time period. When incorporating this convex hull into a multi-period formulation and including it in large-scale energy system models, the improved LP relaxations can better prevent simultaneous charging and discharging, and the tighter MIP could positively affect the solving time. We demonstrate this with illustrative case studies of a unit commitment problem and a transmission expansion planning problem.
@@ I-A Motivation
To investigate the optimal investment and operation decisions of storage, we need to formulate realistic models for devices capable of storing energy, such as pumped hydro storage, electric vehicles [srilakshmiEnergyRegulationEV2022], thermal storage, and some forms of demand response [singabhattuDistributedEnergyResources2017]. In some cases, we also want to incorporate reserves, which enable more flexibility by representing the possible increase/decrease to the scheduled amount of charged/discharged energy[morales-espanaHiddenPowerSystem2017]. Mixed-Integer Programming (MIP) allows us to correctly model the optimal operation of storage, using both binary and continuous variables. Binary variables are required to correctly model reserves and to ensure that the storage cannot charge and discharge simultaneously [xuFactoringCycleAging2018, goAssessingEconomicValue2016]. Constraints in MIP models can describe the characteristics and capabilities of storage units, such as their minimum and maximum state-of-charge, and their (dis)charging capacity.
However, there are two big problems that arise when using these storage models to solve large-scale problems. The first is that computational problems arise when MIP storage models are included in large-scale energy system optimization models. MIP models generally do not scale well, since existing algorithms for finding the optimal solution to MIP models run in exponential time (in the worst case). A common practice to reduce the computation time of these large energy models is to use a decomposition technique, or to relax the integrality restrictions in the MIP formulations. In either case, the Linear Programming (LP) relaxation of the storage model is solved, which can be solved much faster in practice. However, the problem with these LP models for storage operation is that their solutions can enforce simultaneous charging and discharging, although this cannot be executed in practice. To address both of these challenges, we improve the LP approximation by tightening the MIP models for optimal storage investment and operation, including reserves. When incorporated in a large-scale model, these tighter MIP formulations could speed up the solving time. When incorporated into a large-scale LP model, the improved LP relaxations can better prevent simultaneous charging and discharging.
@@ I-B Background & Literature Review
Charging and discharging at the same time essentially increases losses, which commonly increases costs[arroyoUseConvexModel2020]. Therefore, the optimal solution to the LP relaxation usually does not favor this, but it is beneficial in some cases, such as in the event of negative prices. However, simultaneous charging and discharging can usually not be executed in practice. There are exceptions, such as composite battery systems[elsaadanyLinearModelAggregated2025] or other specific cases with multiple and/or fast storage units. Here, simultaneous charging and discharging might actually be executable by charging a fraction of the time period or storage units, and discharging another fraction[shenModelingArbitrageEnergy2020]. Still, this is not executable when the storage unit is already at maximum capacity, and it is typically not possible in practice at all for generic storage units. Therefore, solutions in which this occurs are unwanted and can lead to large errors.
@@
Some examples of situations in which simultaneous charging and discharging in the model solutions can lead to large errors are the following. Arroyo et al. [arroyoUseConvexModel2020] present a transmission expansion planning case study, where the optimal investment plan is actually not feasible in practice, because simultaneous charging and discharging is needed to execute it. Furthermore, immediate load recovery cannot be guaranteed in demand response models with load shifting when simultaneous up and down shifting of load is allowed [morales-espanaClassifyingModellingDemand2022]. Lastly, significant errors can occur when decomposition algorithms are used to solve large MIP problems faster [goAssessingEconomicValue2016]. In these decomposition frameworks, the LP relaxation (of the operations/second stage) is solved many times, thus allowing simultaneous charging and discharging and resulting in errors.
@@
Several attempts to prevent simultaneous charging and discharging in convex problems have been made, such as including pre-contingency operating costs in the objective function [wenEnhancedSecurityConstrainedUnit2016] or adding conditions for the roundtrip efficiency [zhaoUsingElectricalEnergy2018]. However, Arroyo et al. [arroyoUseConvexModel2020] pointed out that it can still occur and illustrated this in two case studies: a unit commitment problem and a transmission expansion planning problem. Thus, an improved LP-approximation of storage models that can better prevent simultaneous charging and discharging would be a big step toward more useful and better scalable storage models.
@@
An improved LP relaxation can also speed up the solving time of MIP storage models. The computational performance of MIP models is highly dependent on their tightness and their size (in terms of the number of constraints and variables)[morales-espanaTightCompactMILP2013]. We say that an MIP model is tight if its LP relaxation describes the convex hull of its solutions. We say that MIP model $A$ is tighter than another MIP model $B$ if the LP relaxation of MIP model $A$ is contained inside the LP relaxation of MIP model $B$, and it is strictly smaller than the other (so $\text{LP}(A)\subsetneq\text{LP}(B)$). Typically, the tightness of MIP models can be improved by adding cuts (additional constraints). However, this also increases the size of the model, which in turn could negatively impact the computational performance.
@@
In related fields, such as unit commitment modeling, much research has been done on finding tighter MIP models[rajanMinimumPolytopesUnit2005, morales-espanaTightCompactMILP2013, gentileTightMIPFormulation2017, huangCuttingPlanesSecurityconstrained2021]. These works take into account the important trade-off between model tightness and model size. For example, Damci-Kurt et al.[damci-kurtPolyhedralStudyProduction2016] first consider the unit commitment problem with ramping constraints for a single generator in just two time periods. They give a complete linear description of the convex hull of its feasible solutions. This formulation can be incorporated into a multi-generator and multi-period formulation, resulting in a formulation that is tighter than the previously used ramping models, but of similar size. They then consider the multi-period case, and find exponential classes of facet-defining inequalities. Though these improve the tightness of the multi-period problem formulation, they heavily increase the size of the model. It is only beneficial for the solving time if these constraints are added to the formulation using a separation algorithm, when needed, instead of a priori.
@@
Tighter formulations for specific storage devices have been found, such as by Baldick et al.[baldickOptimizationFormulationsStorage2023] for pumped storage hydroelectric generators. For a basic storage operation MIP model, Pozo [pozoConvexHullFormulations2023] recently obtained two different convex hull formulations: a vertex representation and a hyperplane representation. Though not clearly stated in their paper, their formulations describe the convex hull of the problem for only one time period. Additional constraints would be needed to describe the convex hull for a multi-period problem, resulting in a model that likely grows exponentially in size w.r.t. the number of time periods. However, the single-period convex hull formulation can be incorporated into a multi-period formulation, which is then tighter than the commonly used formulation. They demonstrate that incorporating this multi-period LP formulation in a multi-period Set-Point Tracking problem prevents simultaneous charging and discharging more often than the LP relaxation of the commonly used MIP model. Yu et al.[yuTightPowerEnergy2023] presented the same findings, as well as some additional facets for three periods.
@@
However, there are several downsides to the LP formulations obtained by Pozo[pozoConvexHullFormulations2023] and Yu et al.[yuTightPowerEnergy2023]. First, we observe that the formulation by Pozo contains several redundant constraints. Half of the constraints can be removed, which decreases the size of the model. Second, the convex hull by Pozo formulation does not contain the original binary variable. Therefore, their multi-period LP approximation cannot be used to solve the original multi-period MIP storage operation problem. Last, neither Pozo nor Yu et al. consider reserves or investment decisions in their models.
@@ I-C Contributions
In summary, the main contributions of this paper are the following: 1) We provide a convex hull formulation for optimal storage operation for one period, which consists of fewer constraints than the formulation found by Pozo[pozoConvexHullFormulations2023]. We also provide the corresponding tight MIP model and include it in a multi-period formulation, which can be used to solve the original multi-period storage problem, which is not possible with the LP model by Pozo[pozoConvexHullFormulations2023]. 2) We provide tighter MIP models for extended problems that include reserves and investment decisions. We prove that their LP-relaxations provide the convex hull of the feasible solutions to the MIP models for one time period, which guarantees that no tighter MIP or better LP approximation for one time period exists. 3) When incorporated in any type of multi-period energy investment/operation system model, we illustrate that the LP relaxations of the tighter MIP formulations can decrease the frequency of occurrences of simultaneous charging and discharging.
The improved MIP formulations and LP relaxations can be used for many different types of planning problems that include storage, as well as for transmission lines. When solving a large-scale MIP problem, tightness of each of the components of the model is an important factor in the solving time. The tighter MIP storage formulations could positively affect the solving time, as shown by Yu et al. [yuTightPowerEnergy2023]. Furthermore, the improved LP approximation can reduce errors when using decomposition frameworks, as explained earlier.
@@ II-A Modeling Storage Operation Problems
Let us consider a storage unit with $\underline{E}=$ 5MWh, $\overline{E}=$50MWh, $\overline{P^{\mathrm{C}}}=\overline{P^{\mathrm{D}}}=$10MW, and $\eta^{\mathrm{C}}=\eta^{\mathrm{D}}=$ 0.5 (for simple computational purposes). Now suppose the unit is already at maximum capacity in a certain time period $t^{*}-1$, so $e_{t^{*}-1}=50$. In the event of negative prices, or some situation where the surrounding energy system needs to get rid of energy, simultaneous charging and discharging might be beneficial, as explained in the introduction. The BO-LP model allows this, so it may output an optimal solutions such as $(e_{t^{*}-1},p^{\mathrm{C}}_{t^{*}},p^{\mathrm{D}}_{t^{*}},\delta_{t^{*}})=(50,8,2,0.8)$ for the next time period $t^{*}$. The state of charge at the end of period $t^{*}$ is then $e_{t^{*}}=e_{t^{*}-1}+\eta^{\mathrm{C}}p^{\mathrm{C}}_{t^{*}}\Delta-\frac{1}{\eta^{\mathrm{D}}}p^{\mathrm{D}}_{t^{*}}\Delta=50+0.5\cdot 8\cdot 1-2\cdot 2\cdot 1$ =50+44=50=50+4-4=50. So the model allows a solution where it charges as much as it discharges, keeping the battery at the same state of charge, and satisfying all the constraints. However, this can never be executed in practice. Thus, we want to find and LP model that prevents this.
@@ II-B Modeling Storage Operation Problems Including Reserves
As explained earlier, storage can be used to balance the power grid, which is particularly challenging when more renewable energy systems are implemented. Operating reserves, specifically spinning reserves, are useful in this scenario [morales-espanaHiddenPowerSystem2017]. The up/down reserve provided by a unit is generally defined as the available capacity that the unit can output on top of its scheduled amount in each time period [morales-espanaMIPFormulationJoint2014]. For storage units specifically, this means that if the storage unit is scheduled to discharge energy, up reserves can be provided by discharging more energy than scheduled, and down reserves by discharging less (or even charging energy). When charging, up reserves can be provided by charging less (or even discharging), and down reserves by charging more.
@@
$e_{t},\ p^{\mathrm{C}}_{t},\ p^{\mathrm{D}}_{t},\ r^{\mathrm{C}+}_{t},\ r^{\mathrm{C}-}_{t},\ r^{\mathrm{D}+}_{t},\ r^{\mathrm{D}-}_{t}\in\mathbb{R}_{\geq 0}\quad\hskip-5.0pt$ $\forall t\in T$ (2i)
@@
A common way to model storage operation reserves with an MIP formulation is by adapting the constraints from the BO-MIP model, resulting in the Basic Operation incl. Reserves MIP model (BOR-MIP model)[liuSecuredReserveScheduling2021]. Here, (2a) and (2b) replace (1b), since the reserves affect the bounds of the energy storage level and vice versa. Constraint (2c) replaces (1c), bounding the amount of energy that can be charged. Constraint (2c) also bounds the down reserves $r^{\mathrm{C}-}_{t}$ that can be realized by charging more than planned, and (2e) bounds the up reserves $r^{\mathrm{C}+}_{t}$, which can be realized by charging less energy than the planned $p^{\mathrm{C}}_{t}$. Similarly, (2d) replaces (1d), and together with (2f) imposes bounds on $p^{\mathrm{D}}_{t}$, $r^{\mathrm{D}+}_{t}$ and $r^{\mathrm{D}-}_{t}$. Decision variables $r^{-}_{t}$ and $r^{+}_{t}$ then represent the total amount of up and down reserves that can be provided in time period $t$. These total reserves are obtained in (2g) and (2h) and bounded by some value $R^{+}$ and $R^{-}$, respectively.
In this paper, we obtain the convex hull of the BOR-MIP model for one time period. We want to point out that there is an issue with modeling reserves using the BOR-MIP model, namely that it does not fully exploit the flexibility of a storage unit. It only allows limited reserves. We explain this further in Appendix -C. Here we also point out issues with the formulation presented by Momber et al.[momberPEVStorageMultiBus2014]. This formulation adapts the constraints such that the model does fully exploit the flexibility of the storage unit, but it might not ensure that there is enough capacity to provide the reserves. For those still interested in using the formulation by Momber et al., we provide tight formulations in the online companion[elgersmaOnlineCompanionTight2024, Section 4].